A particular reactant decomposes with a half‑life of 111s when its initial concentration is 0.366M.

The same reactant decomposes with a half‑life of 231s when its initial concentration is 0.176M.

Determine the reaction order.

What is the value and units of the rate constant for this reaction?
1
Expert's answer
2020-03-02T06:56:44-0500

To know the order of reactions, one must know the equations of the reactions rate constants.

  1. For the first-order reaction: k1=0.693/t1/2 , where t1/2 - half-life. Rate constants of the same reaction has to be the same, but when its initial concentration is 0.366M then k1=0.0062 , and when its initial concentration is 0.176M then k1=0.003. In conclusion, this is't first-order reaction.
  2. For the second-order reactions: k2=1/(t1/2*C0) , where C0 - initial concentration.

k2=1/(111*0.366)=0.0246 M -1 s -1

k2=1/(231*0.176)=0.0246 M -1 s -1

So, this reaction of decomposition of a particular reactant is a second-order reaction, and its rate constant is: k2=0.0246 M -1 s -1